A vessel of 58L capacity is well insulated and is divided equally by a rigid conducting

diaphragm. Initially both halves contain air at pressure of 137.8 Kpa and 413.4 Kpa

and temperature of 27C and 177C respectively. Find the increase of entropy of the

system


1
Expert's answer
2022-03-03T13:03:27-0500

"\Delta S=\frac Q{\Delta T}=\frac{V(p_2-p_1)}{T_2-T_1}=106.6~\frac JK."